# PiFlow: Principle-aware Scientific Discovery with Multi-Agent Collaboration

### v2

@@ Appendix E Distinction from Advanced Prompt Engineering
Appendix F Proof of the Theorem
We recall the elements here and formally proof the convergence along with boundary of the system. Here we denote $\pi$ is the language model policy from policy space $\Pi$, $f^{*}$ is the acquisition function from function space $\mathcal{F}$, $h_{t}$ is the hypothesis at time step $t$, $H_{t-1}=\{h_{1},h_{2},\dots,h_{t-1}\}$ is the history of hypotheses, $v^{*}$ is the SQ achievable by any hypothesis, and $I(h_{t};f^{*}\mid H_{t-1})$ is the conditional mutual information.
@@ Appendix F Proof of the Theorem
Since the information theory proofed that, the variance of a function can be bounded by mutual information, akin to entropy bounds $H(f^{*})\leq\log(|\mathcal{F}|)$, for the first term, we have
@@
Appendix G Algorithmic Realization of the Min-Max Framework
G.1 Practical Proxies for Regret and Information Gain
@@
Approximating exploration (maximizing information gain). The second term in our objective, maximizing the mutual information $I(h_{t};f^{*}|H_{t-1})$, encourages the selection of hypotheses that are most informative about the underlying scientific landscape. A direct computation of mutual information is often intractable. Consequently, Algorithm 1 employs a practical proxy via the compute_exploration_scores function. This is achieved by analyzing the semantic diversity of principles using their sentence level embeddings (e.g., obtained from QwenMax model). We posit that principles which are semantically distant from those already tested are more likely to reveal novel information about the function $f^{*}$. Therefore, we use the cosine distance between a candidate principle’s embedding and the embeddings of previously explored principles as a heuristic for information gain. High exploration scores, $S_{exploration}\in\mathbb{R}^{|\mathcal{T}_{t}|}$, are assigned to conceptually novel principles. This heuristic is theoretically grounded, as detailed in Appendix G.2.
@@
G.2 The Rational of Semantic Distance as a Proxy for Information Gain
@@
With these premises, we can construct the logical argument. The mutual information term $I(h_{t};f^{*}|H_{t-1})$ quantifies the expected reduction in uncertainty about $f^{*}$ after observing the outcome of hypothesis $h_{t}$, given the history of observations $H_{t-1}$. As established in our convergence proof (Section F), this information gain is related to the posterior variance of the outcome of $h_{t}$:
@@
Appendix H Illustrative Example: Application to Nanohelix Optimization
@@
Let’s assume the goal is to find the nanohelix geometry (described at Appendix S.1) with the maximum g-factor:
@@
Appendix I Analysis of Baselines Response
@@
Backend these outputs, the performance difference stems from the level of dynamic and strategic guidance. For example, MPO’s reasoning is local and tactical. It generates a fixed "how-to" checklist. This plan dictates the operational steps but is agnostic to the underlying scientific principles driving the experiments, and cannot reason about why its plan is failing, leading to hypotheses that yield lower outcomes. The Vanilla agent here uses the exact same powerful LLM (QwenMax) and has access to the same experimental tools. Its poor performance demonstrates that even a capable LLM, without strategic guidance, explores aimlessly. The dramatic performance gap between Vanilla and PiFlow (e.g., AUC improving from 35.96% to 63.51% in the NHO task) is direct evidence that our contributions, i.e., (a) structuring knowledge into an explicit principle space, and (b) applying a rigorous Min-Max optimization strategy to navigate it.
@@
Appendix J Performance Comparison with Bayesian Optimization
@@
Appendix K Plug-and-Play Integration with ChemToolAgent
We integrate PiFlow with ChemToolAgent (Yu et al., 2024) on the Molecular Bio-activity Optimization (MBO) task to demonstrate the generalization of PiFlow. The integration follows a Plug-and-Play design, where PiFlow acts as a strategic guidance layer, as shown in Figure 5. ChemToolAgent serves as both the Hypothesizer and Experimenter, leveraging its built-in tools (such as searching molecules’ formula from PubMed and website) to propose molecular hypotheses, while PiFlow’s outer-loop guidance provided Explore and Refine commands to ensure systematic and efficient discovery dynamically. As shown in Table 6, this collaborative process successfully evolved its strategy from basic principles to a high-value molecular design (pChEMBL of 5.90) over 8 iterations without any architecture-level modifications over ChemToolAgent, demonstrating a clear synergy.
@@ 3 Explore ChemToolAgent generated a principle $p_{3}$ and proposed an invalid formula null
Appendix L Theoretical Computational Complexity
@@
Appendix M Cost-Effectiveness Analysis
@@
Appendix N Ablation: Temporal Dynamics of Principle Evaluation
@@
Appendix O Ablation: The Impact of Foundation Models
Influence of Foundation Models. To evaluate the impact of different models, we replace LLMs of $\mathcal{A}_{H}$ and $\mathcal{A}_{P}$ evaluation. The $\mathcal{A}_{E}$ is responsible for tool usage, consistently used QwenMax to ensure functional tool interaction. We evaluate several state-of-the-art LLMs (e.g., Claude-3.7-sonnet (Anthropic, 2025), GPT4.1-mini (OpenAI, 2025), Gemini-2.5-pro-exp-0325 (DeepMind, 2025), Qwen3-32B and QwenMax (Yang et al., 2024b)) on the Nanohelix Optimization (NHO) task with three times of random seeds as initialization, and the results are summarized in Table 9.
@@
Among these models, QwenMax demonstrates the highest AUC at 63.51% and a strong SQ of 76.82%. Claude-3.7-sonnet achieves the highest SQ at 78.50% with an AUC of 38.60%. Other models like GPT4.1-mini (AUC 41.68%, SQ 66.38%) and Qwen3-32B (AUC 37.51%, SQ 58.76%) also show competent performance. Gemini-2.5-pro-exp-03-25 yields an AUC of 28.43% and an SQ of 69.64%. This suggest that, with the different system prompts of Hypothesis Agent and Experiment Agent, the selection of the model may affect the overall performance.
In summary, the performance variation across different models highlights the importance of model selection for specific tasks within the PiFlow-enhanced MAS.
@@
Appendix P Ablation: The Impact of Initial Principle Quality
@@
Takeaway: This study showcases the defining feature of PiFlow: strategic robustness. Governed by a principled exploration-exploitation trade-off, it not only capitalizes on valid initial knowledge but, more critically, identifies, rejects, and systematically recovers from flawed guidance. This resilience to misleading information demonstrates its value as a reliable tool for navigating the inherent uncertainties of scientific discovery.
Appendix Q Ablation: Sensitivity to Hyperparameter $\lambda$
Q.1 Experiment Settings
@@
  • For broad, novel, or theoretically uncertain domains. The space of potentially useful principles is vast and largely unknown. In these cases, one should use a larger lambdalambda value (e.g., 0.7-0.9 or more) to prioritize Exploration. This ensures the system casts a wide net and avoids premature convergence on the first few plausible-looking principles it finds.
@@
  • For well-defined, mature, or theoretically constrained domains. The space of effective principles is likely smaller and more focused. Here, a smaller lambdalambda value (e.g., 0.1-0.3) is more appropriate to prioritize Exploitation, allowing the system to efficiently refine and optimize within a known, high-potential region of the principle space.
@@
As a direction for future work, we are exploring a dynamic lambdalambda scheduling policy. Such a policy would start with a high lambdalambda to encourage initial exploration and automatically decrease it as the system identifies promising regions, thus transitioning smoothly from exploration to exploitation without manual intervention.
@@
Appendix R Benchmark Formulation and Rationale
@@
R.1 Surrogate Models as Validation Functions
@@
R.2 The AUC Metric in High-Variance Tasks
@@
Appendix S Experimental Setup
S.1 Task 1: Nanohelix Optimization (NHO)
S.1.1 Problem Statement
@@
S.1.3 Development
Figure 12: The $r^{2}$ of the surrogate model for g-factor prediction.
@@
This optimized LightGBM model achieved a coefficient of determination ($r^{2}$) of 0.9802, as shown in Figure 12, indicating that the model explains 98.02% of the variance in the g-factor prediction given any structural parameters. This high level of accuracy enables reliable exploration of the parameter space without requiring computationally expensive DFT simulations for each parameter combination.
S.2 Task 2: Bio-activity Optimization (MBO)
S.2.1 Problem Statement
@@
Figure 13: The $r^{2}$ of the surrogate model for bio-activity prediction.
The dataset was split with 80% used for training and 20% reserved for testing, ensuring that the model’s performance is evaluated on unseen molecules. The final model achieved a coefficient of determination ($r^{2}$) of 0.910 on the test set, as shown in Figure 13, demonstrating strong predictive capability across diverse molecular structures.
@@
S.3 Task 3: Superconductor Critical Temperature Optimization (SPO)
S.3.1 Problem Statement
@@
S.3.2 Development
Figure 14: The $r^{2}$ of the surrogate model for $T_{c}$ prediction.
@@
We developed a Multi-Layer Perceptron (MLP) model to predict $T_{c}$. The model was trained on the training set and evaluated on the test set. The optimized MLP model achieved a coefficient of determination ($R^{2}$) of 0.9133 on the test set, as shown in Figure 14[cite: 1]. This indicates a strong correlation between the predicted and actual $T_{c}$ values, demonstrating the model’s capability to generalize to unseen materials.

### v3

@@ Appendix B Distinction from Advanced Prompt Engineering
Appendix C Proof of the Theorem
We recall the elements here and formally prove the convergence along with boundary of the system. Here we denote $\pi$ is the language model policy from policy space $\Pi$, $f^{*}$ is the acquisition function from function space $\mathcal{F}$, $h_{t}$ is the hypothesis at time step $t$, $H_{t-1}=\{h_{1},h_{2},\dots,h_{t-1}\}$ is the history of hypotheses, $v^{*}$ is the SQ achievable by any hypothesis, and $I(h_{t};f^{*}\mid H_{t-1})$ is the conditional mutual information.
@@ Appendix C Proof of the Theorem
Since the information theory proved that, the variance of a function can be bounded by mutual information, akin to entropy bounds $H(f^{*})\leq\log(|\mathcal{F}|)$, for the first term, we have
@@
Appendix D Algorithmic Realization of the Min-Max Framework
D.1 Practical Proxies for Regret and Information Gain
@@
Approximating exploration (maximizing information gain). The second term in our objective, maximizing the mutual information $I(h_{t};f^{*}|H_{t-1})$, encourages the selection of hypotheses that are most informative about the underlying scientific landscape. A direct computation of mutual information is often intractable. Consequently, Algorithm 1 employs a practical proxy via the compute_exploration_scores function. This is achieved by computing the semantic distance of principles using their sentence level embeddings (e.g., obtained from QwenMax model). We posit that principles which are semantically distant from those already tested are more likely to reveal novel information about the function $f^{*}$. Therefore, we use the cosine distance between a candidate principle’s embedding and the embeddings of previously explored principles as a heuristic for information gain. High exploration scores, $S_{exploration}\in\mathbb{R}^{|\mathcal{T}_{t}|}$, are assigned to conceptually novel principles. This heuristic is theoretically grounded, as detailed in Appendix D.2.
@@
Historical information management.
We manage historical information via two independent parts: (a) agent memory for MAS to iterate the hypothesis-testing and (b) persistent principle pool for PiFlow layer to guide the process.
  1. 1.
    For MAS part, agent history differs per agent. As the MAS follows round-robin order, there is the risk of context collapse. To prevent this, each agent maintains an independent MessageBuffer, retaining the most recent $M=10$ messages. This ensures agents focus on the immediate logical flow without being overwhelmed by noise.
  2. 2.
    For PiFlow, as a plugin layer, it has its own memory. Crucially, our principle-aware design decouples reasoning history from scientific insight. While chat logs may be truncated, generated principles and validation scores represent compressed, high-value knowledge stored in a separate, persistent global pool. As per Algorithm 1 (Lines 14-23), even if an agent forgets a specific past conversation, PiFlow retrieves the refined principles from the global pool to guide the next step. This ensures the logic chain remains intact throughout the discovery process.
With the design of decoupling, PiFlow only need to collect historical principles, hypotheses and outcomes to optimize the strategy (explore, refine or validate which temporary principle), and steers the learning process by giving guidance to MAS.
Realworld efficiency (stopping by condition).
In practical deployment, PiFlow’s Min-Max optimization naturally drives the system toward convergence, enabling precise threshold-based stopping: In the Molecular Bio-activity Optimization (MBO) task, PiFlow-MAS improved pChEMBL from 2.69 to 6.44 within just 4 effective iterations, and surpassed the target (6.5) to reach 7.65 by the 6th iteration. This demonstrates that with a practical stop condition, PiFlow achieves success using only  25% of the allocated benchmark budget. As detailed in Section 5.4, our Min-Max framework provides a theoretical guarantee that prevents infinite loops of low-quality exploration, ensuring resource efficiency even in the intensive tasks noted by the Reviewer.
D.2 The Rational of Semantic Distance as a Proxy for Information Gain
@@
With these premises, we can construct the logical argument. The mutual information term $I(h_{t};f^{*}|H_{t-1})$ quantifies the expected reduction in uncertainty about $f^{*}$ after observing the outcome of hypothesis $h_{t}$, given the history of observations $H_{t-1}$. As established in our convergence proof (Section C), this information gain is related to the posterior variance of the outcome of $h_{t}$:
@@
Appendix E Illustrative Example: Application to Nanohelix Optimization
@@
Let’s assume the goal is to find the nanohelix geometry (described at Appendix P.1) with the maximum g-factor:
@@
Appendix F Analysis of Baselines Response
F.1 Response Analysis of Baselines
@@
Backend these outputs, the performance difference stems from the level of dynamic and strategic guidance. For example, the reasoning of MPO and two AI Scientsit methods are local and tactical. MPO generates a fixed "how-to" checklist. This plan dictates the operational steps but is agnostic to the underlying scientific principles driving the experiments, and cannot reason about why its plan is failing, leading to hypotheses that yield lower outcomes. While The AI Scientist v2 and AI-Researcher logically select the “node” or candidate, their lack of hypothesis diversity often leads to premature convergence and performance stagnation. Additionally, the Vanilla agent uses the exact same powerful LLM (QwenMax) and has access to the same experimental tools. Its poor performance demonstrates that even a capable LLM, without strategic guidance, explores aimlessly. The dramatic performance gap between Vanilla and PiFlow (e.g., AUC improving from 35.96% to 63.51% in the NHO task) is direct evidence that our contributions, i.e., (a) structuring knowledge into an explicit principle space, and (b) applying a rigorous Min-Max optimization strategy to navigate it.
@@
F.2 Statistical Significant Test on Metrics
We compared PiFlow against AI-Researcher (Tang et al., 2025) (NeurIPS 2025) and The AI Scientist v2 (Yamada et al., 2025). As shown in Table 5, PiFlow’s targeted uncertainty reduction (Min-Max) proves superior to generic autonomous loops.
Table 5: Performance comparison between PiFlow, The AI Scientist v2, and AI-Researcher.
Task Method SQ (%) P-value (SQ) AUC (%) P-value (AUC)
NHO PiFlow 76.82 $\pm$ 4.54 63.51 $\pm$ 11.18
The AI Scientist v2 56.67 $\pm$ 5.79 0.0103 49.27 $\pm$ 1.84 0.1547
AI-Researcher 53.12 $\pm$ 1.06 0.0091 46.45 $\pm$ 2.42 0.1122
MBO PiFlow 84.55 $\pm$ 29.63 64.57 $\pm$ 23.65
The AI Scientist v2 62.47 $\pm$ 8.02 0.3250 36.32 $\pm$ 10.62 0.1630
AI-Researcher 95.66 $\pm$ 7.45 0.5870 15.56 $\pm$ 12.71 0.2625
SPO PiFlow 34.85 $\pm$ 1.19 21.51 $\pm$ 2.80
The AI Scientist v2 29.85 $\pm$ 2.68 0.0663 28.43 $\pm$ 3.13 0.0470
AI-Researcher 25.69 $\pm$ 3.23 0.0277 16.36 $\pm$ 2.72 0.0842
The results show that, PiFlow achieves statistically significant improvement (p-value <0.05<0.05) on NHO and SPO tasks. On the MBO task, PiFlow remains competitive with higher efficiency (AUC 64.57% vs 42.72%), though SQ variance is higher. This confirms the advantage of PiFlow in complex, uncertainty-heavy environments.
@@ F.2 Statistical Significant Test on Metrics
Appendix G Performance Comparison with Bayesian Optimization
@@
In summary, PiFlow addresses the challenge of semantic hypothesis generation, where BO, Generic Algorithms and Reinforcement Learning methods face significant barriers in representation. PiFlow operates in an open-ended semantic space (natural language principles), allowing it to be Plug-and-Play without domain-specific encoding.
@@
Appendix H Plug-and-Play Integration with ChemToolAgent
We integrate PiFlow with ChemToolAgent (Yu et al., 2024) on the Molecular Bio-activity Optimization (MBO) task to demonstrate the generalization of PiFlow.
The integration follows a Plug-and-Play design, where PiFlow acts as a strategic guidance layer, as shown in Figure 5. ChemToolAgent serves as both the Hypothesizer and Experimenter, leveraging its built-in tools (such as searching molecules’ formula from PubMed and website) to propose molecular hypotheses, while PiFlow’s outer-loop guidance provided Explore and Refine commands to ensure systematic and efficient discovery dynamically. As shown in Table 7, this collaborative process successfully evolved its strategy from basic principles to a high-value molecular design (pChEMBL of 5.90) over 8 iterations without any architecture-level modifications over ChemToolAgent, demonstrating a clear synergy.
@@ 3 Explore ChemToolAgent generated a principle $p_{3}$ and proposed an invalid formula null
Appendix I Theoretical Computational Complexity
@@
The decision complexity of $O(t^{2}\cdot d)$ is not a bottleneck.
We profiled the runtime in the NHO task. Over 24 iterations, the entire PiFlow framework accounted for only 9.4% of the total PiFlow-MAS runtime, while the surrogate model took 0.2%. We also quantitatively profiled the optimization efficiency in the NHO task. The result confirms that the algorithmic complexity of PiFlow translates into superior search efficiency: For baselines, Vanilla Agent System achieves SQ=50% taking 2416s; The other baseline MPO achieves SQ=52% taking 506s and baseline ReAct even fails to reach such level of SQ in all its 918s. However, our PiFlow achieves SQ=52% (similar bar) by only 427s, delivering such a result using less than 1/5 of the Vanilla Agent’s time (a 5.6x speedup), and is 1.2x faster than the baseline MPO.
The results indicate that, the vast majority of time is consumed by LLM inference. The specific similarity calculation ($t^{2}$) takes milliseconds, which is structurally distinct from—and negligible compared to—LLM generation (seconds) or real-world wet-lab experiments (hours/days).
@@
Appendix J Cost-Effectiveness Analysis
@@
Appendix K Ablation: Temporal Dynamics of Principle Evaluation
@@
Appendix L Ablation: The Impact of Foundation Models
Influence of Foundation Models. To evaluate the impact of different models, we replace LLMs of $\mathcal{A}_{H}$ and $\mathcal{A}_{P}$ evaluation. The $\mathcal{A}_{E}$ is responsible for tool usage, consistently used QwenMax to ensure functional tool interaction. We evaluate several state-of-the-art LLMs (e.g., Claude-3.7-sonnet (Anthropic, 2025), GPT4.1-mini (OpenAI, 2025), Gemini-2.5-pro-exp-0325 (DeepMind, 2025), Qwen3-32B and QwenMax (Yang and others, 2024)) on the Nanohelix Optimization (NHO) task with three times of random seeds as initialization, and the results are summarized in Table 10.
@@
Among these models, QwenMax demonstrates the highest AUC at 63.51% and a strong SQ of 76.82%. Claude-3.7-sonnet achieves the highest SQ at 78.50% with an AUC of 38.60%. Other models like GPT4.1-mini (AUC 41.68%, SQ 66.38%) and Qwen3-32B (AUC 37.51%, SQ 58.76%) also show competent performance. Gemini-2.5-pro-exp-03-25 yields an AUC of 28.43% and an SQ of 69.64%. These quantitative results reveal that, while efficiency (AUC) is model-dependent—ranging from 38.60% (Claude-3.7) to 63.51% (QwenMax), Solution Quality (SQ) benefits consistently from stronger reasoners.Specifically, Claude-3.7-sonnet achieves the highest SQ of 78.50%, outperforming the efficiency-optimized QwenMax (76.82%) by $\sim$1.7%.
In summary, these results confirm that PiFlow generalizes well, effectively converting superior base model capabilities into higher-quality optimization outcomes.
@@
Appendix M Ablation: The Impact of Initial Principle Quality
@@
Remark (In the case of no pre-defined principles.) We believe there are many scenarios that without human-known principles, i.e., no pre-defined principles that LLMs can propose. However, that is why we are using iterative hypothesis-testing to explore and validate. We discuss this real-world situation below, i.e., if there are no pre-defined principles:
  1. (a)
    Firstly, initialization for solving the Cold Start. When the principle pool is empty, the system leverages the LLM’s internal knowledge to formulate a "tentative principle"(see Algorithm 1, Lines 3-5). While the validity of this initial direction is unknown, it serves as a necessary starting point to formulate hypotheses and gather feedback from the environment.
  2. (b)
    Secondly, self-correction via Min-Max. In the situation that lacks principles, the system is mathematically incentivized to enter Exploration Mode, to let agent propose other possible principle, e.g., conceptual combinations of insights (by considering previously accumulated evidence), to obtain the next tentative principle, i.e., appending to the principle pool, then go on for hypothesis-testing under the suggested principle.
Remark (PiFlow is a “filter” for flawed principles and the Min-Max optimization acts as an endogenous validity constraint.) With the objective in Eq. 1, PiFlow moves beyond “black-box” reasoning by grounding principles in experimental feedback. It does not require an external filter because the correction mechanism is mathematically embedded in the objective:
  1. (a)
    On one hand, we have theoretical guarantee. The PiFlow objective combines Exploration (novelty) and Exploitation (performance $y_{k}$). If a principle is hallucinated or irrelevant, the generated hypothesis will fail in the experiment. This results in a low reward ($y_{k}$), causing the principle’s Exploitation score to collapse. Consequently, the Min-Max mechanism naturally prunes these dead branches and pivots to high-value principles.
  2. (b)
    On the other hand, we have empirical proof. We provide concrete evidence of this robustness in Appendix M. Instead of random hallucinations, we deliberately initialize PiFlow with Expert-Incorrect principles (misleading physics rules). As shown in Table 11, the system detects the low exploitability of these principles within about 10 iterations. PiFlow successfully abandons the flawed priors and recovers to converge on high-quality solutions (about 70% SQ).
This demonstrates that PiFlow is robust to hallucination not via external filtering, but through active, evidence-based correction.
Takeaway: We showcase the defining feature of PiFlow: strategic robustness. Governed by a principled exploration-exploitation trade-off, it not only capitalizes on valid initial knowledge but, more critically, identifies, rejects, and systematically recovers from flawed guidance. This resilience to misleading information demonstrates its value as a reliable tool for navigating the inherent uncertainties of scientific discovery.
Appendix N Ablation: Hyperparameters Sensitivity and Surrogate Noise
The weight of exploitation and exploration ($\lambda$).
@@
  • For broad, novel, or theoretically uncertain domains. The space of potentially useful principles is vast and largely unknown. In these cases, one should use a larger $\lambda$ value (e.g., 0.7-0.9 or more) to prioritize Exploration. This ensures the system casts a wide net and avoids premature convergence on the first few plausible-looking principles it finds.
@@
  • For well-defined, mature, or theoretically constrained domains. The space of effective principles is likely smaller and more focused. Here, a smaller $\lambda$ value (e.g., 0.1-0.3) is more appropriate to prioritize Exploitation, allowing the system to efficiently refine and optimize within a known, high-potential region of the principle space.
@@
As a direction for future work, we are exploring a dynamic $\lambda$ scheduling policy. Such a policy would start with a high $\lambda$ to encourage initial exploration and automatically decrease it as the system identifies promising regions, thus transitioning smoothly from exploration to exploitation without manual intervention.
Analysis of principle set size ($|P|$).
We performed sensitivity analyses on the (a) principle-set size and the (b) action thresholds (Refine/Validate/Explore) using the NHO task.
Table 13: Ablation study on initial principle set size.
Configuration SQ (%) AUC (%)
init $|P|=3$ 76.82 $\pm$ 4.54 63.51 $\pm$ 11.18
init $|P|=6$ 72.47 $\pm$ 9.34 49.341 $\pm$ 7.17
init $|P|=9$ 76.13 $\pm$ 5.55 47.19 $\pm$ 4.10
As shown in Table 13, varying $|P|\in\{3,6,9\}$ showed stable performance. While smaller sets ($|P|=3$, SQ=76.8%) focus search efficiently, larger sets ($|P|=9$, SQ=76.1%) maintain high quality despite increased exploration costs.
Analysis of action thresholds.
We tested Strict mode (Refine > 0.8, Validate > 0.5), Default (Refine > 0.7, Validate > 0.4), and Loose mode (Refine > 0.6, Validate > 0.3) settings.
As shown in Table 14, PiFlow is robust across a reasonable range. The “Default” and “Loose” settings perform similarly well. Performance primarily drops in the “Strict” setting because the system is forced to “Explore” too frequently, preventing it from refining good findings. This confirms the thresholds are not brittle magic numbers.
Table 14: Comparison of different threshold configurations.
Configuration SQ (%) AUC (%)
Loose (0.6/0.3) 70.65 $\pm$ 9.66 52.18 $\pm$ 13.68
Default (0.7/0.4) 76.82 $\pm$ 4.54 63.51 $\pm$ 11.18
Strict (0.8/0.5) 62.35 $\pm$ 11.48 46.85 $\pm$ 12.24
Robustness against surrogate noise (addressing reward hacking).
In our experiments, we uses a surrogate model to serve as the validation end for hypothesis-testing. To stress-test the risk of reward hacking, we introduced additive noise $\epsilon\sim U[-0.2,0.2]$ (approx. 10% of the effective range) to the surrogate model of NHO task. Other parameters are kept, as with the same configuration in Table 1.
Table 15: Robustness analysis of PiFlow compared to baselines under noise.
Method Noise Level SQ (%) AUC (%)
PiFlow 0% (Clean) 76.82 $\pm$ 4.54 63.51 $\pm$ 11.18
PiFlow 10% 73.65 $\pm$ 2.54 61.82 $\pm$ 1.95
MPO (Baseline) 10% 50.46 $\pm$ 5.04 41.74 $\pm$ 4.87
Vanilla (Baseline) 10% 51.38 $\pm$ 2.88 42.64 $\pm$ 3.09
ReAct (Baseline) 10% 50.00 $\pm$ 5.05 40.92 $\pm$ 4.12
The results are shown in Table 15, as baselines collapsed (chasing noisy gradients), PiFlow maintained high SQ. This indicate that PiFlow demonstrates superior robustness, validating that scientific principles act as effective regularizers against noise.
A case study of literature retrieval integration.
We conducted a case study integrating a Search Agent (using Semantic Scholar API, returning top-5 paper abstract only) into the NHO task:
  1. (a)
    The 1st retrieval provides one key variable. The Search Agent retrieved papers linking "two-turn SiO2 nanohelices" to "circular dichroism control (directly related to our objective value)". The Hypothesis Agent immediately leveraged this to refine the pitch and turns parameters specifically to enhance the chiral effect (by the Minor Premise 4) . This led to a rapid g-factor jump from 0.52 -> 1.58. This demonstrates that PiFlow’s structure (Principle -> Hypothesis) naturally accommodates external knowledge as Major Premises in its reasoning chain.
  2. (b)
    The 2nd retrieval provides theoretical support. When the progress come to the study of number of turns, Search Agent finds literatures about “two-turn SiO2 nanohelixes” and circular dichroism: “The circular dichroism described with g-factor is also presented here as a function of wavelength… leads to flexible control over the circular dichroism”. Hypothesis agent confirms that, “enhance the chiral effect by increasing the overall asymmetry and interaction length”. While the outcome is incremental, from 1.585 to 1.590, this searching ensures the adjustements is appropriate.
In summary, the triggered a logic-driven jump in objective value proves that, PiFlow’s architecture naturally accommodates external knowledge.
@@
Appendix O Benchmark Formulation and Rationale
@@
O.1 Surrogate Models as Validation Functions
@@
O.2 The AUC Metric in High-Variance Tasks
@@
Appendix P Experimental Setup
P.1 Task 1: Nanohelix Optimization (NHO)
P.1.1 Problem Statement
@@
P.1.3 Development
@@
This optimized LightGBM model achieved a coefficient of determination ($r^{2}$) of 0.9802, indicating that the model explains 98.02% of the variance in the g-factor prediction given any structural parameters. This high level of accuracy enables reliable exploration of the parameter space without requiring computationally expensive DFT simulations for each parameter combination.
P.2 Task 2: Bio-activity Optimization (MBO)
P.2.1 Problem Statement
@@
The dataset was split with 80% used for training and 20% reserved for testing, ensuring that the model’s performance is evaluated on unseen molecules. The final model achieved a coefficient of determination ($r^{2}$) of 0.910 on the test set, demonstrating strong predictive capability across diverse molecular structures.
@@
P.3 Task 3: Superconductor Critical Temperature Optimization (SPO)
P.3.1 Problem Statement
@@
P.3.2 Development
@@
We developed a Multi-Layer Perceptron (MLP) model to predict $T_{c}$. The model was trained on the training set and evaluated on the test set. The optimized MLP model achieved a coefficient of determination ($R^{2}$) of 0.9133 on the test set. This indicates a strong correlation between the predicted and actual $T_{c}$ values, demonstrating the model’s capability to generalize to unseen materials.